OpenStudy (love_333):

For the following two questions find the volume of the square pyramids shown. 1: A. 484cm^3 B. 44cm^3 C. 528cm^3 D. 792cm^3 2: A. 4,000ft^3 B. 6,000ft^3 C. 6,250ft^3 D. 7,500ft^3 Medal + fan for both answers!

OpenStudy (love_333):

#1

OpenStudy (love_333):

#2

OpenStudy (love_333):

@Mehek14 Any Ideas on this?

OpenStudy (erak):

For the first one, it's just A = lwh/3

OpenStudy (erak):

For the second one you have to visualize the diagonal line to be the hypotenuse of a triangle that forms with the base and tip of the pyramid. Since it's a square based pyramid, you know the base is 30 ft all around, and exactly in the middle would be 15 ft. Therefore C^2 = A^2 + B^2 25^2 = 15^2 + B^2 B^2 = 625 - 225 B^2 = 400 B = 20 ft <-- this is your height of the pyramid Now use A = lwh/3 again.

OpenStudy (love_333):

528! Thx :) On number two I guess I did it wrong? I got 3,000 and thats not a choice...

OpenStudy (erak):

Your dimensions would be 30x30x20

OpenStudy (love_333):

hmm.. I'm not sure what I did wrong the first time around, but I got 6,000 that time! Thank you!